OpenStudy (anonymous):

log3^x + log3^5 = 2 how can i solve this?

sam (.sam.):

log3^x + log3^5 = 2 \[\log_3(5x) = 2\] Use \(log_a(b)=c ~~~~~~<-->~~~~~~a^c=b\) \[3^2=5x\] Solve

OpenStudy (anonymous):

so do i multiply 3 by 5

sam (.sam.):

\[3^2=5x\] Do the exponent first, then divide by 5 \[9=5x\] \[x=\frac{9}{5}\]